A 56 0-kg person holding two 0.750-kg bricks stand
Answer #1

At all times the total momentum must be conserved. Since the person was initially at rest, so the initial momentum is zero. Hence the final momentum after the person throws the bricks must also be zero.

\small (m_{person}+m_{skateboard})\times v_{recoil}+ m_{bricks}\times v_{bricks}=0 \\\Rightarrow v_{recoil}= -\frac{m_{bricks}\times v_{bricks}}{m_{person}+m_{skateboard}}=-\frac{2\times 0.75 \times 18}{56+3}=-0.46 ms^{-1}

So the recoil velocity is 0.46 m/s and the negative sign implies that the person moves in a direction opposite to that of the bricks.